# InsightBench: Evaluating Business Analytics Agents Through Multi-Step Insight Generation

### v1

@@ Abstract
Data analytics is essential for extracting valuable insights from data that can assist organizations in making effective decisions. We introduce InsightBench, a benchmark dataset with three key features. First, it consists of 31 datasets representing diverse business use cases such as finance and incident management, each accompanied by a carefully curated set of insights planted in the datasets. Second, unlike existing benchmarks focusing on answering single queries, InsightBench evaluates agents based on their ability to perform end-to-end data analytics, including formulating questions, interpreting answers, and generating a summary of insights and actionable steps. Third, we conducted comprehensive quality assurance to ensure that each dataset in the benchmark had clear goals and included relevant and meaningful questions and analysis. Furthermore, we implement a two-way evaluation mechanism using LLaMA-3-Eval as an effective, open-source evaluator method to assess agents’ ability to extract insights. We also propose AgentPoirot, our baseline data analysis agent capable of performing end-to-end data analytics. Our evaluation on InsightBench shows that AgentPoirot outperforms existing approaches (such as Pandas Agent) that focus on resolving single queries. We also compare the performance of open- and closed-source LLMs and various evaluation strategies. Overall, this benchmark serves as a testbed to motivate further development in comprehensive data analytics and can be accessed here: https://github.com/ServiceNow/insight-bench.
Figure 1: Existing benchmarks (top) assess the agents’ ability to solve single data analytics tasks with defined outputs. InsightBench (bottom) specializes in evaluating the effectiveness of LLM-based agents for comprehensive data analytics processes. This includes evaluating the agents’ ability to recommend the best specific tasks to address a specific goal, extracting insights across descriptive, diagnostic, predictive, and prescriptive categories, and summarizing both findings and recommended next steps (as demonstrated by AgentPoirot).
@@ 1 Introduction
Businesses frequently leverage vast datasets to perform data analytics to uncover insights, discover patterns, and analyze trends in order to support effective decision-making (McAfee and Brynjolfsson, 2012; Colson, 2019; Bean, 2022). The task of end-to-end data analysis starts with stating a high-level goal; the analyst then alternates between identifying key questions to explore, and extracting valuable insights from their answers, working towards the goal. This iterative process continues until they have a comprehensive summary of insights and recommended actions (shown in Figure 1).
Analysts often carry out such tasks by leveraging tools like Jupyter notebooks and dashboards (Bean, 2022; Yin et al., 2023), which, unfortunately, require both significant human effort as well as data science and domain expertise. Fortunately, agents based on large language models (LLMs) have emerged as promising assistants to users in performing data science tasks (OpenAI, 2022, 2023). But these agents (such as Pandas Agent (LangChain, 2024) and Code Interpreter (OpenAI, 2024)) only focus on solving narrow, single-step code-completion (or data analytics) tasks (E.g., “What is the R2 value for a linear regression model on this dataset?”).
@@
This inspired us to propose InsightBench, which consists of 31 datasets whose structures have been acquired from the ServiceNow (ServiceNow, 2024) platform, which focuses on business operations and workflows. This platform allows us to generate synthetic tabular data that mimics real-life enterprise data. We used it to create datasets for Incident, User, Finance, Inventory, and Enterprise Goal Management. We chose this data because an agent capable of effectively extracting insights from it can be deployed to help organizations optimize their business operations.
@@ 2 InsightBench – An Enterprise Data Analysis Benchmark
InsightBench’s data were derived from the ServiceNow (ServiceNow, 2024) platform, which is used to manage business workflows crucial for enterprise operations. It consists of tables that store and manage records and data entities relevant to various organizational functions. Common tables in ServiceNow include incident tables (for incident records), user tables (for user information), and asset management tables (to oversee infrastructure). InsightBench consists 31 tabular datasets acquired from the ServiceNow platform covering five distinct themes (see Figure˜3).
Overview of Benchmark Creation.    InsightBench is an automated data analysis benchmark that can rigorously assess the performance of LLM-based data analysis agents in real-world scenarios. Building InsightBench consisted of four stages: 1) selecting relevant data tables from the ServiceNow data tables and extracting a list of relevant columns to define the schema (Section 2.1.1), 2) formulating trends to inject in the data (Section 2.1.2), 3) creating synthetic data entries to populate the tables that follow the trend formulated in the previous step (Section 2.1.3), and 4) creating a ground-truth analysis notebook for the generated data (Section 2.1.4). See Figure 2 for an overview of the multi-stage process. A detailed list of the datasets sorted by themes is presented in Section˜B.4. We propose a multi-step evaluation mechanism to measure the performance of an agents on InsightBench (See Section 2.2)
@@ 2.1.2 Planting Insights
We embed systematic anomalies and trends into the datasets to certain types of columns in the datasets, referred to as controllable columns. These trends were implemented using standard mathematical models or algorithms, such as regression, to ensure consistency and detectability of the planted insights. For instance, we manipulated the time-to-resolution (TTR) trend for incidents in a service management dataset to reflect an increasing trend over time. This was achieved by using a linear model to generate the TTR based on the creation date of each incident: $\text{TTR}=1+\text{slope}\cdot(\text{incident\_open\_date}-\text{data\_start\_date})$, where (slope) is a parameter that controls the rate of increase in resolution time (See Figure 2(b)).
@@ 2.1.4 Ground-Truth Analysis Notebooks
A key component of constructing InsightBench was developing 31 expert-annotated Jupyter notebooks, each tailored for a specific context, such as incident management or financial operations. The notebook structure and contents are as follows (see Figure 2(c)):
@@ 2.2 Evaluating Agents on InsightBench
To evaluate agents on InsightBench, we compare the ground-truth annotations ($GT$) in Jupyter notebooks against the list of insights ($A$) provided by an agent. Formally speaking, we compare two natural language texts. While multiple metrics have been proposed in the past, like ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTSCore (Zhang et al., 2019), they do not align well with human preferences. Recently, G-Eval (Liu et al., 2023) was shown to be highly aligned with human preferences for a variety of tasks, including text summarization. Since our task involves comparing two free-form sentences, we adopt an LLM-based evaluator as well. Specifically, we use LLaMA-3-Eval, a variant of G-Eval that uses LLaMA-3-70b instead of a GPT model. We use LLaMA-3-Eval as the primary metric to measure the correctness of agent-provided insights. We chose LLaMA-3-Eval over G-Eval because LLaMA-3-70b is open-sourced, allowing us to a) avoid API costs for using GPT models from OpenAI and b) fix the model weights to obtain a stable and reliable evaluation metric, unlike G-Eval, whose output scores can change due to periodic updates to GPT endpoints.
@@
We average the summary-level and insight-level scores for all 31 datasets in InsightBench to obtain a measure of the agent’s performance. Additionally, we compute scores using ROUGE-1 for comparison.
@@ 2.3 Dataset Statistics
InsightBench covers a distinct range of business analytics themes and varying difficulty levels, ensuring a comprehensive evaluation of the agents. The benchmark includes 31 datasets spread across five key topics. These topics were chosen to reflect typical scenarios encountered in enterprise settings, each requiring clear and well-defined insights. The distribution across the topic is shown in Figure 3, and the data tables used are outlined in Section˜B.3
Figure 3: This diagram illustrates the scope of InsightBench, categorizing 31 distinct datasets into five key thematic areas.
Distribution by Difficulty. The datasets are assigned a difficulty level on a scale from 1 (easy) to 4 (hard). Easy (Level 1-2) comprises 9 datasets that primarily involve direct data retrieval and basic analysis. Medium (Level 3) includes 12 datasets that may require the integration of multiple data sources or applying moderate data transformations. Hard (Level 4-5) consists of 10 datasets that require the calculation of multiple intermediate quantities or significant transformation of data variables. A detailed list of datasets with their respective topics and difficulty range is outlined in Section˜B.4
Distribution of Insight Types: Overall, there are 45 Descriptive insights, 40 Diagnostic insights, 15 Predictive insights, and 10 Prescriptive insights in InsightBench. Discovering these insights will allow the agents to not only uncover and understand the data but also forecast trends and recommend precise, actionable strategies.
@@ 2.4 Quality Assurance
To ensure our datasets are of high quality, we designed a questionnaire and an interface that displays our dataset to users. We had 20 volunteers with basic data science skills evaluate whether our datasets were accurately described, contained relevant and interesting questions, and provided substantial insights. The exact questions used, the interface design, and the results are shown in B.5. The average rating was 4 out of 5 across these questions, and we received helpful comments that improved parts of the dataset.
@@ 3 Experimental Setup
Effect of Question Diversity. We restrict the agent to generate only “descriptive" questions as follow-ups instead of letting it generate different types of questions.
@@ 4 Benchmark Results
PA (gpt-4o) $3±\pm0$.01 0.52 $\pm$0.01 0.34 $\pm$0.01 0.38 $\pm$0.02
Ours (gpt-4o) 0.31 $\pm$0.01 0.59 $\pm$0.01 0.36 $\pm$0.08 0.41 $\pm$0.02
     - w/ generic goal 0.29 $\pm$0.01 0.39 $\pm$0.02 0.29 $\pm$0.06 0.31 $\pm$0.13
Ours (gpt-4-turbo) 0.29 $\pm$0.00 0.55 $\pm$0.01 0.34 $\pm$0.08 0.32 $\pm$0.03
     - w/ generic goal 0.27 $\pm$0.00 0.37 $\pm$0.01 0.27 $\pm$0.02 0.26 $\pm$0.09
Ours (gpt-3.5-turbo) 0.32 $\pm$0.01 0.48 $\pm$0.02 0.26 $\pm$0.12 0.30 $\pm$0.06
     - w/ generic goal 0.29 $\pm$0.01 0.35 $\pm$0.02 0.23 $\pm$0.02 0.24 $\pm$0.06
Ours (llama-3-70b) 0.31 $\pm$0.01 0.51 $\pm$0.03 0.35 $\pm$0.01 0.32 $\pm$0.01
     - non-diverse follow-ups 0.28 $\pm$0.00 0.50 $\pm$0.02 0.31 $\pm$0.01 0.27 $\pm$0.02
     - w/ generic goal 0.26 $\pm$0.02 0.34 $\pm$0.01 0.22 $\pm$0.01 0.21 $\pm$0.02
@@
Table 1 also shows the performance of AgentPoirot that uses a generic goal. We notice a drastic overall decrease in the ROUGE-1 and LLaMA-3-Eval scores for all the backbones. For instance, the insight-level LLaMA-3-Eval score for AgentPoirot (gpt-4o) drops from 0.59 ($\pm$ 0.01) to 0.39 ($\pm$ 0.02) and the summary-level LLaMA-3-Eval score drops from 0.41 ($\pm$0.02) to 0.31 ($\pm$0.13) (see “Ours (gpt-4o)” v/s “Ours (gpt-4o) w/ generic goal" rows in Table 1). This highlights the importance of including a SMART goal in the agent’s performance.
@@
Table 1 includes the performance of AgentPoirot when generating only a single type of follow-up questions (see “Ours (llama-3-70b) non-diverse follow-ups"). We observe a notable decrease in ROUGE-1 scores and a slight decrease in LLaMA-3-Eval scores, confirming that discovering a diverse set of questions during data analysis leads to better performance.
@@
The increase in volume of incidents is seen only for one particular category i.e. Hardware The trend in assigned incidents varies significantly across different assignment groups, with the “Hardware" group having the highest number of incidents (314), followed by “Software" (56), “Network" (48), “Database" (43), and “Service Desk" (39). The category ‘Hardware’ has shown the highest increase in the number of incidents over the given period, with a total increase of 32.0 incidents.
@@ 5 Related Work
Data Science Benchmarks    With enhanced abilities in code generation, utilization of LLM-based data analytics assistants is becoming increasingly prevalent. This has led to the development of numerous data science benchmarks (Lai et al., 2022; Chandel et al., 2022; Zan et al., 2022; Hu et al., 2024; Zhang et al., 2024b). DS-1000 (Lai et al., 2022) focuses on code generation for diverse data science questions sourced from StackOverflow, emphasizing functional correctness in solutions. Similarly, InfiAgent-DABench (Hu et al., 2024) assesses the end-to-end problem-solving ability of LLMs by requiring them to interact with an execution environment to answer questions based on provided CSV files. DSP (Chandel et al., 2022), evaluates a model’s proficiency through a code-infilling task within Jupyter Notebooks. DSEval (Zhang et al., 2024b) focuses on the overall behavior of data science agents without getting into the nuances of code generation techniques. On the other hand, Text2SQL and tabular data reasoning benchmarks assess the ability of models to parse queries, extract information, and synthesize the retrieved data to formulate responses (Katsogiannis-Meimarakis and Koutrika, 2021; Zhong et al., 2018; Chen et al., 2021b). While existing benchmarks primarily evaluate agents on single-step code generation, InsightBench shifts the focus to multi-step analysis(Delen and Ram, 2018).
@@
LLM Evaluation Frameworks    Current LLM evaluation frameworks primarily focus on measuring the competence of LLM-based agents in handling structured outputs Zhang et al. (2023). Existing evaluations predominantly rely on pre-formatted prompts to assess code completion (Wu et al., 2023; Zhang et al., 2024a; Yao et al., 2023). While recent advancements have seen autonomous agents specializing in intricate data science tasks, including analysis, visualization, and modeling (Qian et al., 2023, 2024; Zhang et al., 2023), evaluations for these methods often depend on extensive human effort (Cheng et al., 2023) or use more powerful LLMs to assess the output (Dubois et al., 2023; Belyi et al., 2024). G-Eval (Liu et al., 2023) is a recent technique used to evaluate the quality of free-form texts in terms of factuality and coherence. In this work, we use a variation of G-Eval to score how well the predicted insights are aligned with the ground-truth insight.
@@
Text-to-Analytics Agents    Chen et al. (2023) explore the application of GPT variants within a data visualization context, highlighting the strengths and limitations of these models. More recent LLM-based data analysis agents include Code Interpreter (OpenAI, 2024) and Pandas Agent (LangChain, 2024) that are capable of processing multiple data formats and answering questions about them. Ma et al. (2023) propose InsightPilot, an advanced automated tool that leverages LLMs to enhance data exploration by automatically identifying goals and generating targeted intentional queries. Vacareanu et al. (2024) showed that LLMs can also perform regression tasks, enhancing their predictive abilities. Additional studies assess the data analysis capabilities of GPT-4 and propose an end-to-end framework for automating data processes (Cheng et al., 2023; Wang et al., 2023, 2024; Hong et al., 2024). Inspired by InsightPilot, we propose AgentPoirot that can perform end-to-end data analysis that includes extracting descriptive, diagnostic, predictive, and prescriptive insights.

### v2

@@ Abstract
Data analytics is essential for extracting valuable insights from data that can assist organizations in making effective decisions. We introduce InsightBench, a benchmark dataset with three key features. First, it consists of 100 datasets representing diverse business use cases such as finance and incident management, each accompanied by a carefully curated set of insights planted in the datasets. Second, unlike existing benchmarks focusing on answering single queries, InsightBench evaluates agents based on their ability to perform end-to-end data analytics, including formulating questions, interpreting answers, and generating a summary of insights and actionable steps. Third, we conducted comprehensive quality assurance to ensure that each dataset in the benchmark had clear goals and included relevant and meaningful questions and analysis. Furthermore, we implement a two-way evaluation mechanism using LLaMA-3 as an effective, open-source evaluator to assess agents’ ability to extract insights. We also propose AgentPoirot, our baseline data analysis agent capable of performing end-to-end data analytics. Our evaluation on InsightBench shows that AgentPoirot outperforms existing approaches (such as Pandas Agent) that focus on resolving single queries. We also compare the performance of open- and closed-source LLMs and various evaluation strategies. Overall, this benchmark serves as a testbed to motivate further development in comprehensive automated data analytics  [^1] Preprint. Under Review..
@@ 1 Introduction
Businesses frequently leverage vast datasets to perform data analytics to uncover insights, discover patterns, and analyze trends in order to support effective decision-making (McAfee & Brynjolfsson, 2012; Colson, 2019; Bean, 2022). The task of end-to-end data analysis starts with stating a high-level goal; the analyst then alternates between identifying key questions to explore, and extracting valuable insights from their answers, working towards the goal. This iterative process continues until they have a comprehensive summary of insights and recommended actions (shown in Figure 1).
Figure 1: Existing benchmarks (top) assess the agents’ ability to solve a highly specific data analytics query with pre-defined output templates. They often require an expert user to ask the questions. InsightBench (bottom), on the other hand, evaluates the LLM-based agents on the complete comprehensive data analytics processes. This includes evaluating the agents’ ability to answer high-level questions by a general user, recommend the best specific tasks to address a specific goal, extracting insights across descriptive, diagnostic, predictive, and prescriptive categories, and summarizing both findings and recommending next steps (as demonstrated by AgentPoirot).
Analysts often carry out such tasks by leveraging tools like Jupyter notebooks and dashboards (Bean, 2022; Yin et al., 2023), which, unfortunately, require both significant human effort as well as data science and domain expertise. Fortunately, agents based on large language models (LLMs) have emerged as promising assistants to users in performing data science tasks (OpenAI, 2022; 2023). But these agents (such as Pandas Agent (LangChain, 2024) and Code Interpreter (OpenAI, 2024)) only focus on solving narrow, single-step code-completion (or data analytics) tasks (E.g., “What is the R2 value for a linear regression model on this dataset?”).
@@
This inspired us to propose InsightBench, which consists of 100 datasets whose structures have been acquired from the ServiceNow (ServiceNow, 2024) platform, which focuses on business operations and workflows. This platform allows us to generate synthetic tabular data that mimics real-life enterprise data. We used it to create datasets for Incident, User, Finance, Inventory, and Enterprise Goal Management. We chose this data because an agent capable of effectively extracting insights from it can be deployed to help organizations optimize their business operations.
@@ 2 InsightBench – An Enterprise Data Analysis Benchmark
InsightBench’s data were derived from the ServiceNow (ServiceNow, 2024) platform, which is used to manage business workflows crucial for enterprise operations. It consists of tables that store and manage records and data entities relevant to various organizational functions. Common tables in ServiceNow include incident tables (for incident records), user tables (for user information), and asset management tables (to oversee infrastructure). InsightBench consists of 100 tabular datasets acquired from the ServiceNow platform covering five distinct themes (see Figure 3).
Overview of Benchmark Creation.    InsightBench is an automated data analysis benchmark that can rigorously assess the performance of LLM-based data analysis agents in real-world scenarios. Building InsightBench consisted of four stages: 1) selecting relevant data tables from the ServiceNow data tables and extracting a list of relevant columns to define the schema (Section 2.1.1), 2) formulating trends to inject in the data (Section 2.1.2), 3) creating synthetic data entries to populate the tables that follow the trend formulated in the previous step (Section 2.1.3), and 4) creating a ground-truth analysis notebook for the generated data (Section 2.1.4). See Figure 2 for an overview of the multi-stage process. A detailed list of the datasets sorted by themes is presented in Section B.3. We propose a multi-step evaluation mechanism to measure the performance of an agents on InsightBench (See Section 2.2)
@@ 2.1.2 Planting Insights
[…]
@@ 2.1.4 Ground-Truth Analysis Notebooks
A key component of constructing InsightBench was developing 100 expert-annotated Jupyter notebooks, each tailored for a specific context, such as incident management or financial operations. The notebook structure and contents are as follows (see Figure 2(c)):
@@ 2.2 Evaluating Agents on InsightBench
To evaluate agents on InsightBench, we compare the ground-truth annotations ($GT$) in Jupyter notebooks against the list of insights ($A$) provided by an agent. Formally speaking, we compare two natural language texts. While multiple metrics have been proposed in the past, like ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), and BERTSCore (Zhang et al., 2019), they do not align well with human preferences. Recently, G-Eval (Liu et al., 2023) was shown to be highly aligned with human preferences for a variety of tasks, including text summarization. Since our task involves comparing two free-form sentences, we adopt an LLM-based evaluator as well. Specifically, we use LLaMA-3-Eval, a variant of G-Eval that uses LLaMA-3-70b instead of a GPT model. We use LLaMA-3-Eval as the primary metric to measure the correctness of agent-provided insights. We chose LLaMA-3-Eval over G-Eval because LLaMA-3-70b is open-sourced, allowing us to a) avoid API costs for using GPT models from OpenAI and b) fix the model weights to obtain a stable and reliable evaluation metric, unlike G-Eval, whose output scores can change due to periodic updates to GPT endpoints.
@@
We average the summary-level and insight-level scores for all 100 datasets in InsightBench to obtain a measure of the agent’s performance. Additionally, we compute scores using ROUGE-1 for comparison.
@@ 2.3 Dataset Statistics
InsightBench covers a distinct range of business analytics themes and varying difficulty levels, ensuring a comprehensive evaluation of the agents. The benchmark includes 100 datasets with a total of 475 insights spread across five key topics. These topics were chosen to reflect typical scenarios encountered in enterprise settings, each requiring clear and well-defined insights. The distribution across the topic is shown in Figure 3, and the data tables used are outlined in Section B.2
Figure 3: This diagram illustrates the scope of InsightBench, categorizing 100 distinct datasets into five key thematic areas.
Distribution by Dataset Category. The benchmark has 20 datasets on Incident Management, 20 datasets on Asset Management, 15 datasets on User Management, 15 datasets on Finance Management, 10 datasets on Goal Management, 10 datasets on Asset & User Management combined, and 10 datasets on Finance & User Management.
Distribution by Difficulty. The datasets are assigned a difficulty level on a scale from 1 (easy) to 4 (hard). Easy (Level 1-2) comprises 30 datasets that primarily involve direct data retrieval and basic analysis. Medium (Level 3) includes 36 datasets that may require the integration of multiple data sources or applying moderate data transformations. Hard (Level 4-5) consists of 34 datasets that require the calculation of multiple intermediate quantities or significant transformation of data variables. A detailed list of datasets with their respective topics and difficulty range is outlined in Section B.3
@@ 2.4 Quality Assurance
To ensure our datasets are of high quality, we designed a questionnaire and an interface that displays our dataset to users. We had 20 volunteers with basic data science skills evaluate whether our datasets were accurately described, contained relevant and interesting questions, and provided substantial insights. The exact questions used, the interface design, and the results are shown in B.4. The average rating was 4 out of 5 across these questions, and we received helpful comments that improved parts of the dataset.
@@ 3 Experimental Setup
Effect of Question Diversity. We restrict the agent to generate only “descriptive” questions as follow-ups instead of letting it generate different types of questions.
@@ 3.1 Baselines
In addition to PandasAgent, we also considered the following baselines: Insightpilot (Ma et al., 2023), but neither the code nor the prompts are available; Data-Copilot (Zhang et al., 2023b), but it is limited to Chinese financial data, which isn’t suitable for our task; OpenAgents (Xie et al., 2023), but the public version only allows 10 requests, which isn’t enough for meaningful benchmarking; InfiAgent-Dabench (Hu et al., 2024), but we excluded it due to its extremely poor performance in this benchmark. It handles only simple tasks like “Generate Python code to compute the mean of a list,” but struggles with more complex queries like “What is the distribution of incidents by category in this dataset?” Lastly, we considered using PowerBI  [^3] https://app.powerbi.com/home, but it is a closed-source tool that is not suitable for large-scale benchmarking. We conducted a small-scale evaluation of PowerBI and found that it obtained low-quality insights and generated inconsistent outputs for the same prompt. Our benchmark emphasizes transparency and reproducibility, therefore we focussed on methods that the scientific community can easily inspect and scrutinize, which is not possible with the aforementioned baselines.
@@ 4 Benchmark Results
PA (gpt-4o) 0.35$\pm$0.03 0.54 $\pm$0.01 0.35 $\pm$0.01 0.40 $\pm$0.04
Ours (gpt-4o) 0.32 $\pm$0.02 0.60 $\pm$0.03 0.37 $\pm$0.09 0.44 $\pm$0.03
   - w/ generic goal 0.30 $\pm$0.03 0.40 $\pm$0.03 0.30 $\pm$0.08 0.33 $\pm$0.12
Ours (gpt-4-turbo) 0.30 $\pm$0.02 0.56 $\pm$0.02 0.35 $\pm$0.08 0.35 $\pm$0.04
   - w/ generic goal 0.28 $\pm$0.01 0.38 $\pm$0.03 0.29 $\pm$0.02 0.27 $\pm$0.11
Ours (gpt-3.5-turbo) 0.34 $\pm$0.01 0.50 $\pm$0.02 0.27 $\pm$0.14 0.31 $\pm$0.06
   - w/ generic goal 0.30 $\pm$0.02 0.36 $\pm$0.03 0.24 $\pm$0.03 0.25 $\pm$0.06
Ours (llama-3-70b) 0.33 $\pm$0.02 0.52 $\pm$0.04 0.36 $\pm$0.01 0.33 $\pm$0.01
   - non-diverse follow-ups 0.29 $\pm$0.01 0.51 $\pm$0.03 0.32 $\pm$0.01 0.28 $\pm$0.03
   - w/ generic goal 0.27 $\pm$0.03 0.35 $\pm$0.01 0.23 $\pm$0.03 0.23 $\pm$0.02
@@
Table 1 also shows the performance of AgentPoirot that uses a generic goal. We notice a drastic overall decrease in the ROUGE-1 and LLaMA-3-Eval scores for all the backbones. For instance, the insight-level LLaMA-3-Eval score for AgentPoirot (gpt-4o) drops from 0.59 ($\pm$ 0.01) to 0.39 ($\pm$ 0.02) and the summary-level LLaMA-3-Eval score drops from 0.41 ($\pm$0.02) to 0.31 ($\pm$0.13) (see “Ours (gpt-4o)” v/s “Ours (gpt-4o) w/ generic goal” rows in Table 1). This highlights the importance of including a SMART goal in the agent’s performance.
@@
Table 1 includes the performance of AgentPoirot when generating only a single type of follow-up questions (see “Ours (llama-3-70b) non-diverse follow-ups”). We observe a notable decrease in ROUGE-1 scores and a slight decrease in LLaMA-3-Eval scores, confirming that discovering a diverse set of questions during data analysis leads to better performance.
@@
The increase in volume of incidents is seen only for one particular category i.e. Hardware The trend in assigned incidents varies significantly across different assignment groups, with the “Hardware” group having the highest number of incidents (314), followed by “Software” (56), “Network” (48), “Database” (43), and “Service Desk” (39). The category ‘Hardware’ has shown the highest increase in the number of incidents over the given period, with a total increase of 32.0 incidents.
@@ 5 Related Work
Data Science Benchmarks    With enhanced abilities in code generation, utilization of LLM-based data analytics assistants is becoming increasingly prevalent. This has led to the development of numerous data science benchmarks (Lai et al., 2022; Chandel et al., 2022; Zan et al., 2022; Hu et al., 2024; Zhang et al., 2024b). DS-1000 (Lai et al., 2022) focuses on code generation for diverse data science questions sourced from StackOverflow. Similarly, InfiAgent-DABench (Hu et al., 2024) assesses the end-to-end problem-solving ability of LLMs by asking questions based on provided CSV files. DSP (Chandel et al., 2022), evaluates a model’s proficiency through a code-infilling task within Jupyter Notebooks. DSEval (Zhang et al., 2024b) focuses on the overall behavior of data science agents without getting into the nuances of code generation techniques. On the other hand, Text2SQL and tabular data reasoning benchmarks assess the ability of models to parse queries, extract information, and synthesize the retrieved data to formulate responses (Katsogiannis-Meimarakis & Koutrika, 2021; Zhong et al., 2018; Chen et al., 2021b). While existing benchmarks primarily evaluate agents on single-step code generation, InsightBench shifts the focus to multi-step analysis (Delen & Ram, 2018).
@@
LLM Evaluation Frameworks    Existing LLM evaluations frameworks focus on handling structured outputs and predominantly rely on pre-formatted prompts to assess code completion (Wu et al., 2023; Zhang et al., 2023a; 2024a; Yao et al., 2023). While recent advancements have seen autonomous agents specializing in intricate data science tasks, including analysis, visualization, and modeling (Qian et al., 2023; 2024; Zhang et al., 2023a), evaluations for these methods often depend on extensive human effort (Cheng et al., 2023) or use more powerful LLMs to assess the output (Dubois et al., 2023; Belyi et al., 2024). Another work uses the “Capture the Flag” principle, where insights are planted into a dataset as flags to evaluate whether models can uncover them (Laradji et al., 2023). G-Eval (Liu et al., 2023) is a recent technique used to evaluate the quality of free-form texts in terms of factuality and coherence. In this work, we use a variation of G-Eval to score how well the predicted insights are aligned with the ground-truth insight.
@@
Text-to-Analytics Agents    Chen et al. (2023) explore the application of GPT variants within a data visualization context, highlighting the strengths and limitations of these models. More recent LLM-based data analysis agents include Code Interpreter (OpenAI, 2024) and Pandas Agent (LangChain, 2024) that are capable of processing multiple data formats and answering questions about them. Ma et al. (2023) propose InsightPilot, an advanced automated tool that leverages LLMs to enhance data exploration by automatically identifying goals and generating targeted intentional queries. Vacareanu et al. (2024) showed that LLMs can also perform regression tasks, enhancing their predictive abilities. Additional studies assess the data analysis capabilities of GPT-4 and propose an end-to-end framework for automating data processes (Cheng et al., 2023; Wang et al., 2023; 2024; Hong et al., 2024). Inspired by InsightPilot, we propose AgentPoirot that can perform end-to-end data analysis that includes extracting descriptive, diagnostic, predictive, and prescriptive insights.